OpenStudy (anonymous):

let f(x)={2x-5 if x<=1 {2x+5 if x>1 then find of f(x) at x=1

OpenStudy (anonymous):

what is the instruction for this? :)

OpenStudy (anonymous):

show all necessary steps clearly

OpenStudy (anonymous):

okay, where does x=1, falls is it on x< =1, or on x>1?

OpenStudy (anonymous):

See this is a piecewise function in which function is defined as 2x-5 for values of x <=1 and 2x + 5 for values of x > 1 . Now we have to find f(x) at x=1 which means it lies in 2x- 5 , therefore f( x) = 2x1 - 5 = -3